OpenStudy (anonymous):

for the given functions f ang g find the following. For parts (a)-(d),also find the domain *Given f(x)=1/x and (f/g)(x)=x+1/x^2-x find the function g.

OpenStudy (anonymous):

that was close. just upside down i think

OpenStudy (anonymous):

oops where did it go?

OpenStudy (anonymous):

haha I thought it was upside down too. g(x) = (x - 1) / (x + 1)

OpenStudy (anonymous):

ok should i rewrite the problem

myininaya (myininaya):

\[(f/g)(x)=\frac{x+1}{x^2-x}=\frac{x+1}{x(x-1)}=\frac{1}{x}*\frac{x+1}{x-1}\] \[=\frac{1}{x} \div \frac{x-1}{x+1}=f \div g\]

myininaya (myininaya):

do you see what g is?

OpenStudy (anonymous):

domain of f, all real numbers except 0. range all number except 0

OpenStudy (anonymous):

domain of g all numbers except -1 range all numbers except 1

OpenStudy (anonymous):

ooo ok im trying to figure it out thanks you guyz

OpenStudy (anonymous):

yw from myininaya

myininaya (myininaya):

yes look what i wrote is it obvious what g is?

OpenStudy (anonymous):

o duh yes lol

myininaya (myininaya):

go to sleep satellite